OpenStudy (anonymous):

Find the length of JK. Give your answer in terms of π. 2 cm2 4 cm2 6 cm2 3 cm2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@GodlyMaths

jimthompson5910 (jim_thompson5910):

what is the circumference?

OpenStudy (anonymous):

I have to find it ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

3.14 x 2?

jimthompson5910 (jim_thompson5910):

times the radius as well

OpenStudy (anonymous):

3.14 x 2 x 5?

OpenStudy (anonymous):

31.4?

jimthompson5910 (jim_thompson5910):

now multiply that by 72/360 to get the piece of the perimeter

OpenStudy (anonymous):

6.28

jimthompson5910 (jim_thompson5910):

6.28 is approximately 2pi

jimthompson5910 (jim_thompson5910):

since 2*pi = 2*3.14 = 6.28

OpenStudy (anonymous):

Okay what do i do next ?

jimthompson5910 (jim_thompson5910):

you're done

OpenStudy (anonymous):

The answer is 6cm^2?

OpenStudy (anonymous):

i mean 2cm^2

jimthompson5910 (jim_thompson5910):

no, 2pi sq cm

jimthompson5910 (jim_thompson5910):

it says "Give your answer in terms of π."

OpenStudy (anonymous):

Find < MNO